# Enhancing Amharic-LLaMA: Integrating Task Specific and Generative Datasets

### v1

@@ Abstract
Large language models (LLMs) have received a lot of attention in natural language processing (NLP) research because of their exceptional performance in understanding and generating human languages. However, low-resource languages are left behind due to the unavailability of resources. In this work, we focus on enhancing the LLaMA-2-Amharic model by integrating task-specific and generative datasets to improve language model performance for Amharic. We compile an Amharic instruction fine-tuning dataset and fine-tuned LLaMA-2-Amharic model. The fine-tuned model shows promising results in different NLP tasks. We open-source our dataset creation pipeline, instruction datasets, trained models, and evaluation outputs to promote language-specific studies on these models.
@@ 1 Introduction
Adopting these LLMs to local languages requires the preparation of a quality instruction dataset. Amharic is one of the Semitic languages under the Afroasiatic language family spoken in Ethiopia with more than 57M speakers. There are numerous task-specific datasets for Amharic (Tonja et al., 2023) compared to other Ethiopian languages. This paper focuses on enhancing the Llama-2-Amharic model [^1] https://huggingface.co/iocuydi/llama-2-amharic-3784m with quality datasets that are created by converting existing datasets in English into instruction-based Amharic datasets. Furthermore, we create new instruction datasets following the approach by Wei et al. (2021).
Llama-2-Amharic model was created by pre-training Llama-2 7B model using open-source Amharic and translated corpus. After performing vocabulary expansion and pre-training, the authors fine-tuned the created model by translating English instruction datasets into Amharic using commercial translation tools. In our research, we aim to improve the performance of the Amharic LLaMA model by integrating task-specific and generative datasets as shown in Table 1. The contributions of this paper are as follows:
@@ 2 Related Work
The work done by Dong et al. (2023) explores how LLMs’ generation, natural language understanding, and problem-solving abilities relate to the data they are trained on and its composition. This work suggests that the amount of composition data is more important for these abilities to show in a low-resource scenario.
Training resources are one of the most important factors for adopting these models to specific languages. Techniques like low-rank adaptation (LoRA) (Hu et al., 2021) worked on reducing the number of trainable parameters using decomposition matrices during training. Quantized LoRA (QLoRA) (Dettmers et al., 2023) improves training time by introducing quantization and other numerical data type tricks. Using self-instructed fine-tuning, the work by Wei et al. (2021); Taori et al. (2023); Cui et al. (2023) showed a new approach to align the generation outputs of the generative models through the application of NLP tasks. These tasks are structured around natural language instruction templates, providing a novel means to guide the model’s generation process toward better adherence to task-specific requirements. LLaMA-Adapter (Zhang et al., 2023b) also shows that it is possible to reduce the fine-tuning time of LLaMa-7B by introducing lightweight adapters on top of the model.
Acquiring and preparing a dataset for instruction fine-tuning presents a significant challenge due to the extensive labor and resources required. There are several ways of acquiring instruction data, including manual dataset curation, using generative models (Wang et al., 2022; Taori et al., 2023), or using machine translation instruction data for training LLMs for specific languages (Cui et al., 2023).
@@
To the same approach of the work by Cui et al. (2023), Llama-2 was also adopted for the Amharic (iocuydi, 2024) language. During pre-training, iocuydi (2024) used an open-source Amharic corpus with some translated corpus from English, and for fine-tuning, available English instruction datasets were translated to Amharic using the Google Translate API and SeamlessM4T. Following the increase of the LLaMa vocabulary size from 32k to 51k, and subsequent pre-training with a large Amharic text corpus, they conducted supervised instruction fine-tuning using machine-translated datasets. Then they evaluated their model using the MMLU (Hendrycks et al., 2020) multiple-choice English dataset by translating to Amharic. The model is available without original Amharic evaluations because no instruction-based datasets exist for Amharic.
@@
Table 1: Dataset used for preparing instruction fine-tuning data. Is new = dataset that was collected by us.
@@ 3 Dataset preparation
In this work, we have converted existing NLP task-specific datasets, like sentiment analysis and machine translation, into instruction datasets. We created an instruction template for each task and developed a data creation pipeline that merges each template instruction with appropriate data from a pre-existing dataset. This approach helps us to create instruction datasets from pre-existing NLP task datasets. For the new NLP task, we focused on collecting a new dataset that can be converted into instruction data. We also created new datasets by tweaking existing datasets. Finally, we included an instruction-tuning dataset converted into Amharic language using machine translation systems. Table 1 shows a detailed distribution of instruction task data.
@@ 3.1 Instruction dataset from existing datasets
We also worked with MasakhaNews (Adelani et al., 2023) which is a benchmark dataset for news topic classification covering 16 languages widely spoken in Africa. It provides an evaluation of baseline models from classical machine learning models and fine-tunes several language models.
To test if our model has the ability to identify names from sentences, we modified MasakhaNER (Adelani et al., 2021), which is a dataset for named entity recognition (NER) in ten African languages. For this work, we created questions to extract only personal names, and we plan to include more in our future works.
@@
For tasks like Amharic text summarization, we used XL-Sum (Hasan et al., 2021), a comprehensive and diverse dataset comprising 1M annotated article-summary pairs. The dataset covers 44 languages, ranging from low to high-resource ones. We utilized the Amharic portion of the dataset in two ways. First, we took the text and prepared an instructional dataset to test our model’s ability to summarize the text. We also created a text expansion task where our model takes the shorter sentence and produces a detailed explanation about the text, the inverse of the text summarization task.
Finally, to prepare training, validation, and testing for machine translation, we used the dataset in the works done by Barrault et al. (2019); NLLB Team et al. (2022). Our training dataset is from WMT19 (Barrault et al., 2019) and validation and testing are from NLLB Team et al. (2022).
The Amharic spell correction dataset is designed to assess the effectiveness of models in correcting Amharic spelling errors, covering common misspellings to advance NLP tools for the language. For this task, we leveraged Amharic BBC news texts from xlsum (Hasan et al., 2021). We also leveraged the text augmentation library nlpaug (Ma, 2019). We introduced some random character augmentations, including insertion, substitution, swapping, deletion and word cropping. This augmentation is done randomly and applied to part of the dataset.
After preparing each dataset, we found that the machine translation dataset we have was significantly larger than the other tasks, so we set maximum threshold of 10k instructions randomly for the training split of each dataset. For validation and testing, we only used one template per task, and we did not expand the data sizes.
@@ 3.2 New Custom Datasets
In Amharic, music, stories, and poems represent fascinating cultural artifacts. To facilitate the training and evaluation of models’ capabilities in processing these tasks, we have created three new datasets. The first track we considered is religious music lyrics generation. There are several types of music lyrics we included in this dataset. We collected above 2k Amharic spiritual song lyrics from WikiMezmur  [^2] https://wikimezmur.org/am. Despite the popularity of non-religious music in Ethiopia, finding a freely available source to include this in our data was difficult; hence, our non-religious music data was smaller than the others. To expand this dataset, we split the lyrics into verses and created a new completion task where the input is the first verse and the output is the remaining whole verse.
@@ 3.3 Translated instruction fine-tuning dataset
During Llama model self-instructed fine-tuning (Touvron et al., 2023), instruction datasets like Alpaca (Taori et al., 2023) and dolly (Conover et al., 2023) have been widely used. In the work iocuydi (2024), machine translation systems were used to translate these datasets into Amharic instruction fine-tuning data. This method is used by most papers that try to adopt Llama models for their language, like Cui et al. (2023). For Amharic versions of alpaca and dolly datasets, we used datasets used by Llama-2-Amharic (iocuydi, 2024) training. We explored the effect of training a model by using only our relatively clean and human-verified data alone and in combination with this machine translation data.
@@ 4 Experiments
We followed Chinese LLaMA (Cui et al., 2023) experiments to perform supervised fine-tuning (SFT) on our dataset using different types of the dataset we created. We used codes available on the Chinese-LLaMA-Alpaca [^9] https://github.com/ymcui/Chinese-LLaMA-Alpaca repository. We used 4, A100 GPUs with the default parameters in the repository. All training’s are done for 3 epochs. All models and evaluation codes will be available in our repository. For MT task we also worked on M2M100 (Fan et al., 2021) and NLLB (NLLB Team et al., 2022) models.
During the evaluation of the models, we used gpt-4-0613 for GPT-4. For our LLaMa-based models, we used a fixed generation parameters across the models.
@@
  • Fine-tuning existing Amharic-LLaMA model using our own dataset.
  • Exploring the effect of combining our dataset with existing machine-translated instruction datasets for Amharic.
  • Fine-tuning existing models for only the machine translation dataset using our larger dataset and exploring.
  • Exploring the effect of prompts in existing and available models for Amharic tasks.
  • Exploring how code mixing affects the performance of models.
@@ 4.1 Datasets
The first set of experiments we conducted involved evaluating the base Llama-2-Amharic model (iocuydi, 2024) on our custom test set, which was created from different NLP task datasets. This will provide us with a baseline performance for Amharic tasks. The next set of experiments used different NLP task datasets that were converted into an instruction dataset by our data generation pipeline. We used Llama-2-Amharic model (iocuydi, 2024), which is pre-trained using the Llama-2 model for the Amharic language and performed supervised instruction fine-tuning on the task datasets. This ensures our model only has access to quality datasets that were adopted from verified NLP tasks. Finally, we combined our instruction dataset with the machine-translated instruction datasets. In the different datasets above, we have capped our training dataset to a maximum of 10k data from individual tasks, as shown in Table 1. We kept fixed instruction and data frequency in our validation and test set to avoid any performance variation because of instruction differences. For machine translation experiments, we created additional data that contains 200k data points from Barrault et al. (2019) and NLLB Team et al. (2022).
@@
Tasks GPT-4 LLaMA-2-amharic ours (Task data) ours (combined data)

### v3

# Walia-LLM: Enhancing Amharic-LLaMA by Integrating Task-Specific and Generative Datasets

@@ Abstract
Large language models (LLMs) have received a lot of attention in natural language processing (NLP) research because of their exceptional performance in understanding and generating human languages. However, low-resource languages are left behind due to the unavailability of resources. In this work, we focus on enhancing the LLaMA-2-Amharic model by integrating task-specific and generative datasets to improve language model performance for Amharic. We compile an Amharic instruction fine-tuning dataset and fine-tuned LLaMA-2-Amharic model. The fine-tuned model shows promising results in different NLP tasks. We also explore the effectiveness of translated instruction datasets compared to the dataset we created. Our dataset creation pipeline, along with instruction datasets, trained models, and evaluation outputs, is made publicly available to encourage research in language-specific models.
Walia-LLM: Enhancing Amharic-LLaMA by Integrating Task-Specific and Generative Datasets
Israel Abebe Azime 1, Atnafu Lambebo Tonja 2,3, Tadesse Destaw Belay 2, Mitiku Yohannes Fuge 4, Aman Kassahun Wassie4, Eyasu Shiferaw Jada, Yonas Chanie5, Walelign Tewabe Sewunetie 6, Seid Muhie Yimam 7 Masakhane NLP, Ethio NLP, 1 Saarland University, Germany, , 2 Instituto Politécnico Nacional, Mexico,3 Lelapa AI,4 AIMS 5 Carnegie Mellon University, 6 Debre Markos University, 7Universität Hamburg, Germany
@@ 1 Introduction
Adopting these LLMs to local languages requires the preparation of a quality instruction dataset. Amharic is one of the Semitic languages under the Afroasiatic language family spoken in Ethiopia with more than 57M speakers. There are numerous task-specific datasets for Amharic (Tonja et al., 2023) compared to other Ethiopian languages. This paper focuses on enhancing the Llama-2-Amharic (Andersland, 2024) model with quality datasets that are created by converting existing datasets in English into instruction-based Amharic datasets. Furthermore, we create new instruction datasets following the approach by Wei et al. (2022).
Llama-2-Amharic model by Andersland (2024) was created by pre-training Llama-2 7B model using open-source Amharic and translated corpus. After performing vocabulary expansion and pre-training, Andersland (2024) fine-tuned the created model by translating English instruction datasets into Amharic using commercial translation tools. In our research, we aim to improve the performance of the Amharic LLaMA model by integrating task-specific and generative datasets, as shown in Table 1. The contributions of this paper are as follows:
@@ 2 Related Work
The work by Dong et al. (2023) explores how LLMs’ generation, natural language understanding, and problem-solving abilities relate to the data they are trained on and its composition. This work suggests that the amount of composition data is more important for these abilities to show in a low-resource scenario.
Using self-instructed fine-tuning, the work by Wei et al. (2022); Taori et al. (2023); Cui et al. (2023) showed a new approach to align the generation outputs of the generative models through the application of NLP tasks. These tasks are structured around natural language instruction templates, providing a novel means to guide the model’s generation process toward better adherence to task-specific requirements. LLaMA-Adapter (Zhang et al., 2023b) also shows that it is possible to reduce the fine-tuning time of LLaMA-7B by introducing lightweight adapters on top of the model.
Acquiring and preparing a dataset for instruction fine-tuning presents a significant challenge due to the extensive labor and resources required. There are several ways of acquiring instruction data, including manual dataset creation, using generative models (Wang et al., 2022; Taori et al., 2023), or using machine translation instruction data for training LLMs for specific languages (Cui et al., 2023).
@@
To the same approach of the work by Cui et al. (2023), Llama-2 was also adopted for Amharic (Andersland, 2024). During pre-training, Andersland (2024) used an open-source Amharic corpus with some translated corpus from English, and for fine-tuning, available English instruction datasets were translated to Amharic using the Google Translate API and SeamlessM4T. Following the increase of the LLaMA vocabulary size from 32k to 51k and subsequent pre-training with a large Amharic text corpus, they conducted supervised instruction fine-tuning using machine-translated datasets. Then, they evaluated their model using the MMLU (Hendrycks et al., 2020) multiple-choice English dataset by translating it into Amharic. The model is available without original Amharic evaluations because no instruction-based datasets exist for Amharic.
@@
Table 1: Dataset used for preparing instruction fine-tuning data. Is new = new custom dataset. Details of each data source are explained in Section 3. Figure 1 shows how a dataset is converted to instruction data using our Data processing Pipeline.
@@ 3 Dataset preparation
In this work, we have converted existing NLP task-specific datasets, like sentiment analysis and machine translation, into instruction datasets. We created an instruction template for each task and developed a data creation pipeline (Figure 1) that merges each template instruction with appropriate data from a pre-existing dataset. This pipeline helps us to create instruction datasets from pre-existing NLP task datasets. For the new NLP task, we focused on collecting a new dataset that can be converted into instruction data. We also created new datasets by tweaking existing datasets. Finally, we included an instruction-tuning dataset converted into Amharic language using machine translation systems. Table 1 shows a detailed distribution of instruction task data.
@@ 3.1 Instruction dataset from existing datasets
We also worked with MasakhaNews (Adelani et al., 2023), a benchmark dataset for news topic classification covering 16 languages widely spoken in Africa. It provides an evaluation of baseline models from classical machine learning models and fine-tunes several language models.
To test if our model has the ability to identify names from sentences, we modified MasakhaNER (Adelani et al., 2021), which is a dataset for named entity recognition (NER) in ten African languages. We created questions to extract only personal names for this work, and we plan to include more in our future works.
Figure 1: Data processing Pipeline. The pipeline creates instruction data from existing task datasets, and from generative datasets, we collected. All instructions, input, and output are in Amharic except for the MT case, as shown in the picture. The data source will not be used during training.
@@
For tasks like Amharic text summarization, we used XL-Sum (Hasan et al., 2021), a comprehensive and diverse dataset comprising 1M annotated article-summary pairs. The dataset covers 44 languages, ranging from low to high-resource ones. We utilized the Amharic portion of the dataset in two ways. First, we took the text and prepared an instructional dataset to test our model’s ability to summarize the text. We also created a text expansion task where our model takes the shorter sentence and produces a detailed explanation of the text, the inverse of the text summarization task.
Finally, we used the dataset by Barrault et al. (2019); NLLB Team et al. (2022) to prepare training, validation, and testing for machine translation. Our training dataset is from WMT19 (Barrault et al., 2019), and validation and testing are from NLLB Team et al. (2022).
The Amharic spell correction dataset is designed to assess the effectiveness of models in correcting Amharic spelling errors, covering common misspellings to advance NLP tools for the language. We leveraged Amharic BBC news texts from XL-Sum (Hasan et al., 2021) for this task. We also leveraged the text augmentation library nlpaug (Ma, 2019). We introduced some random character augmentations, including insertion, substitution, swapping, deletion, and word cropping. This augmentation is done randomly and applied to part of the dataset.
After preparing each dataset, we found that the machine translation dataset we have was significantly larger than the other tasks, so we set a maximum threshold of 10k instructions randomly for the training split of each dataset. For validation and testing, we only used one template per task, and we did not expand the data sizes. More dataset examples and explanations are found in the Appendix A.
@@ 3.2 New Custom Datasets
In Amharic, music, stories, and poems represent fascinating cultural artifacts. We have created three new datasets to facilitate the training and evaluation of models’ capabilities in processing these tasks. The first track we considered is religious music lyrics generation. We included several types of music lyrics in this dataset. We collected the above 2k Amharic spiritual song lyrics from WikiMezmur [^1] https://wikimezmur.org/am. Despite the popularity of non-religious music in Ethiopia, finding a freely available source to include this in our data was difficult; hence, our non-religious music data was smaller than the others. To expand this dataset, we split the lyrics into verses and created a new completion task where the input is the first verse and the output is the remaining whole verse.
@@ 3.3 Translated instruction fine-tuning dataset
During Llama model self-instructed fine-tuning (Touvron et al., 2023), instruction datasets like Alpaca (Taori et al., 2023) and dolly (Conover et al., 2023) have been widely used. In the work Andersland (2024), machine translation systems were used to translate these datasets into Amharic instruction fine-tuning data. This method is used by most papers that try to adopt Llama models for their language, like Cui et al. (2023). For Amharic versions of alpaca and dolly datasets, we used datasets used by Llama-2-Amharic (Andersland, 2024) training. We explored the effect of training a model by using only our relatively clean and human-verified data alone and in combination with this machine translation data.
Figure 2: Full training pipeline that summarizes the work done.
@@ 4 Experiments
We followed Chinese LLaMA (Cui et al., 2023) experiments to perform supervised fine-tuning (SFT) on our dataset using different types of the dataset we created. Figure 2 shows the full training pipeline that summarizes the overall experiment steps we followed. We used codes available on the Chinese-LLaMA-Alpaca [^8] https://github.com/ymcui/Chinese-LLaMA-Alpaca repository. We used 4 A100 GPUs with the default parameters in the repository. All training is done for three epochs. All models and evaluation codes will be available in our repository. For the MT task, we also worked on M2M100 (Fan et al., 2021) and NLLB (NLLB Team et al., 2022) models.
During the evaluation of the models, we used gpt-4-0613 for GPT-4. For our LLaMA-based models, we used fixed generation parameters across the models. We also evaluated various models that purportedly support this language but excluded them due to their inability to perform the required tasks. This is further discussed in Appendix A.
Our main experiment includes:
@@
  • Fine-tuning the model using a dataset detailed in Table 1, referred to as Walia (task data). Unlike the approach taken in the Llama-2-Amharic model (Andersland, 2024), this experiment did not incorporate machine-translated instructional data.
  • Fine-tuning the model using Walia (combined data), which consists of our prepared dataset along with the machine-translated instructional data previously utilized in the Llama-2-Amharic model (Andersland, 2024).
  • Fine-tuning Walia MT, the MT model we trained to perform the machine translation task in our dataset. In this experiment, we used only MT datasets from Table 1 and scaled them to 200k data rather than 20k as shown in the table.
  • Exploring the effect of prompts in existing and available models for Amharic tasks. Additionally, how code mixing affects the performance of models. This include experiments discussed in 4.3.
@@ 4.1 Datasets
The first set of experiments we conducted involved evaluating the base Llama-2-Amharic model (Wei et al., 2022) on our custom test set, which was created from different NLP task datasets. This will provide us with a baseline performance for Amharic tasks. The next set of experiments used different NLP task datasets that were converted into an instruction dataset by our data generation pipeline. We used Llama-2-Amharic model (Andersland, 2024), which is pre-trained using the Llama-2 model for the Amharic language and performed supervised instruction fine-tuning on the task datasets. This ensures our model only has access to quality datasets that were adopted from verified NLP tasks. Finally, we combined our instruction dataset with the machine-translated instruction datasets. In the different datasets above, we have capped our training dataset to a maximum of 10k data from individual tasks, as shown in Table 1. We kept fixed instruction and data frequency in our validation and test set to avoid any performance variation because of instruction differences. For machine translation experiments, we created additional data that contains 200k data points from Barrault et al. (2019) and NLLB Team et al. (2022).
4.2 Evaluation Metrics
For generation tasks, we used Rogue (Lin, 2004) scores. We used Rogue scores to evaluate xlm-summarization, reverse summarization, and AmharicQA tasks. We reported Rogue1, Rogue2, and RugueL metrics for generation tasks, but we heavily rely on RogueL for analysis since it focuses on the longest common subsequence rather than n-grams. We observed that most of our generation outputs do not share common n-grams when n is greater than 2, and the generations from systems like GPT-4 tend to be longer where the n-gram comparison methods express the results less. Additionally, we used word-based evaluation metrics, SacreBLEU (Post, 2018) and character-based evaluation metrics, chrF++ (Popović, 2017) automatic evaluation metrics for MT tasks. We added character-based metrics (chrf++) because, for low-resource languages with complex morphologies, chrF++ offers a more robust and adaptable metric compared to word-based metrics like SacreBLEU.
Finally, we performed human evaluation for generative tasks such as music, poetry, and story generation. We sampled 120 individual items and conducted blind reviews using three people for each question. We created a rating system from 1 to 5 with detailed instructions and reported the average rating per task and model.
@@ 4.2 Evaluation Metrics
Figure 4: Scores for machine translation. Amharic to English translation scores (Right) and English to Amharic translation scores( left).
4.3 Prompt based experiments
Throughout our investigation, we observed using only one instruction for a task introduces a high dependency on the prompt, leading to prompt over-fitting. In this case, models fail to do tasks like sentiment classification when presented with different instruction prompts. To deal with this problem, we worked on manually produced templates for each task as shown in Table 1.
Additionally, we experimented with the prompt header part of the dataset shown in 1. The prompt header is additional English text stating, "Below is an instruction that describes a task. Write a response that appropriately completes the request.". Introducing this in models like GPT-4 yielded a significant reduction in classifiable outputs, which highlights the effectiveness of incorporating clear English instructions to steer the model toward the desired outcome.
Tasks GPT-4 LLaMA-2-Am Walia I Walia II
Text summarization 3.34 0.62 1.13 0.80
Text expansion 3.10 3.22 2.05 2.82
Amharic QA 28.23 2.83 5.37 6.25
Table 2: I = Walia (task data), II = Walia (combined data). ROUGEL scores for text summarization, Text expansion and Amharic QA.
5 Results
Below, we discuss the performance of each model we tested by task type and evaluation strategy.
5.1 Classification Results
For classification tasks, we used two metrics. Our models improve Llama-2-Amharic scores as shown in AfriSenti, MasakhaNews, and QA tasks in Figure 3. The other metrics we reported measure how many times the model returns one of the categories. For the AfriSenti classification task, 759 and 52 out of 1999 test data are not in any of the three classes for Llama-2-Amharic and GPT-4, respectively. Our models reduce these unusable results and do not produce unusable outputs. For MasakhaNews 248, 136, 106, and 3, results are unusable for Llama-2-Amharic, Walia (task data), Walia (combined data), and GPT-4, respectively. In MasakhaNews case, GPT-4 tends to take the lead in producing reasonable outputs.
Tasks GPT-4 LLaMA-2-Am Walia I Walia II
Story generation 2.93 1.00 3.60 1.73
Poem completion 2.53 1.46 1.73 2.26
Poem generation 2.13 1.00 2.46 2.00
Religious Lyrics Gen. 2.86 1.46 1.60 1.46
Religious Lyrics Compl. 3.60 1.40 2.13 1.93
Non religious Lyrics Gen. 3.53 1.00 1.60 2.06
Table 3: I = Walia (task data), II = Walia (combined data). Average blind human evaluation out of 5, for three people in each task. (1) empty or non Amharic text. (2) not written in task format. (3) written in task format but no consistent idea and spelling errors. (4) looks like that specific generation task but has spelling and grammar errors. (5) this looks like a perfect generation of the task. Underlined text indicates cases where we see improvement compared to LLaMA-2-Amharic.
5.2 Generation Results
As explained in Section 4.2, we focus on RogueL metrics for our analysis. Across text summarization and AmharicQA, GPT-4 takes the lead, showing the model’s generation ability is very high. We were able to improve the LLaMA-2-Amharic model’s ability for this task using our data, as shown in Table 2.
We conducted a human evaluation for the models that do not have fixed gold labels, as shown in Table 3. Table 3 result shows that the generation ability of LLlama-2-Amharic can be enhanced by adding generation-specific datasets. Walia lacks an understanding of the specific formatting of texts because of the limitations in our pre-processing. However, it shows significant improvement where the Llama-2-Amharic fails to understand the query.
5.3 Machine Translation (MT)
For the MT task we evaluated two open-source sequence-to-sequence models (M2M100 (Fan et al., 2021) and NLLB (NLLB Team et al., 2022)), GPT-4, Llama-2-Amharic, and our models. Figure 4 shows SacreBLEU and chrF++ results for the above MT models. As shown in the figure, from MT models, GPT-4 showed better results than the other models when using English as the target language. However, our models showed results comparable to the NLLB and m2m100 models and outperformed the Llama-2-Amharic model for the Amharic-English translation. For the English-Amharic translation, the NLLB model outperformed the others in the SacreBLEU score, while our models showed comparable results and outperformed GPT-4, Llama-2-Amharic and m2m100 models in this translation direction. In our MT evaluation, we noticed irregularities between the results of the two evaluation metrics. Since SacreBLEU is a word-based metric, the results show that the scores are too low. This shows that using only automatic evaluation metrics makes interpreting and generalizing the results hard. We will add metrics like human evaluation to evaluate MT results in the future.
6 Conclusion and Future Works
Moving forward, we aim to improve both the quality and volume of the data utilized. Task-specific dataset creations are meant to complement, not replace, language-specific instruction dataset creations, and we plan to work on creating quality instruction datasets in addition to using existing task datasets. We also plan to explore the relevance of using LLMs for evaluation in low-resource languages like Amharic and incorporate LLMs to evaluate our LLMs.
7 Limitations
One limitation we observed in our work is the lack of reliable generation metrics for our tasks. The models tend to generate wordy and explained outputs despite our attempts to design the instruction template specifically. As a solution, we used several metrics that can express one task’s ability and reported the best-suited one.
In our current evaluation of all the models, we observed significant limitations while preforming the spell correction and NER tasks. For Amharic spell correction all four generation models, including GPT-4, try to generate other things related to the text, and the word error rate for all of them is close to 99%.
We have yet to explore the effect of using machine-translated instruction datasets for building language-specific LLMs with regard to introducing cultural bias.
Appendix A Experimental details
We adopted our LLama-2 instruction tuning experimental code from Chinese-LLaMA-Alpaca [^9] https://github.com/ymcui/Chinese-LLaMA-Alpaca repository by Cui et al. (2023). we performed our instruction frinetuning for three epochs on 4 A100 GPUs using the parameters in Table 5. In the generation phase, we used parameters in Table 5 for all models except GPT-4.
Parameter Value
epoch 3
lr 1e-4
lora_rank 8
lora_alpha 32
lora_dropout 0.05
per_device_train_batch_size 1
per_device_eval_batch_size 1
gradient_accumulation_steps 8
Table 4: Training Parameters
Parameter Value
seed 42
do_sample True
min_length None
top_p 1.0
temperature 1.0
top_k 5
repetition_penalty 5.0
length_penalty 1
Table 5: Configuration settings for token generation.
Appendix B Dataset details
Figure 5 shows how we repurposed existing sentiment analysis data to convert it into an instruction dataset. We utilized a task template selected at random from our collection. The number of templates collected for each task is shown in the table. The reason for keeping the prompt header in the English language is discussed in Section 4.3.
To avoid instruction overfitting, as discussed in Section 4.3, we collected a variety of instructions, like the example shown in Figure 6. The example displays different instructions that can be paired with a machine translation dataset to create a machine translation instruction dataset. This step is repeated for all tasks, including tasks like poem generation, which we created by collecting from different websites.
Due to the difficulty we faced in evaluating generation tasks, we employed human evaluation. Figure 7 shows how evaluators scored each generation output for the case of the story generation task.
Figure 5: Example data output from our dataset creation pipeline. 1
Figure 6: Example templates for machine translation task with English translation. By using random instructions for tasks, we ensure that the model does not fit the specific instructions for tasks.
Figure 7: Form used for human annotation with labeling instruction. We see in the figure how one question of the sample story generation task is being validated.
Appendix C Result details
In Table 6, we present detailed scores for text summarizing, text expansion, and Amharic QA. The choice of the right parameter needs further exploration but RogueL is a good metrics to show improvement because it doesn’t depend on specific n-gram similarities.
In Figure 8 we demonstrate that even if the model’s output is not an exact match, we have created a pipeline to verify and identify the correct output from the generated sequence. We limit our-self to GPT-4 as an external model because it’s not explored in other well-known models. Additionally, we reveal that the mala-500 Lin et al. (2024) model produces unrelated outputs, which merits deeper examination.
Figure 8: Example of analysis we did on the model outputs.
Tasks GPT-4 LLaMA-2-Amharic Walia (Task data) Walia (combined data)